OpenStudy (anonymous):

$500 is deposited into an account quarterly for eight years at an interest rate of 5.3% compounded quarterly. How much is in the account at the end of the 8 years?

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(nt) A = 500(1+0.053/4)^(4*8) A = 500(1+0.053/4)^(32) A = 500(1+0.01325)^(32) A = 500(1.01325)^(32) A = 500(1.5238127) A = 761.90635 So 761.91 dollars will be in the account.